### v1

# Visual Cue Enhancement and Dual Low-Rank Adaptation for Efficient Visual Instruction Fine-Tuning

Abstract
Fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity. To address these issues, we propose an efficient fine-tuning framework with two novel approaches: Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). VCE enhances the vision projector by integrating multi-level visual cues, improving the model’s ability to capture fine-grained visual features. Dual-LoRA introduces a dual low-rank structure for instruction tuning, decoupling learning into skill and task spaces to enable precise control and efficient adaptation across diverse tasks. Our method simplifies implementation, enhances visual comprehension, and improves adaptability. Experiments on both downstream tasks and general benchmarks demonstrate the effectiveness of our proposed approach.
1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across a wide range of tasks, showcasing versatility and effectiveness in diverse applications. Existing MLLMs [14, 1, 15] typically integrate pretrained Large Language Models (LLMs) [26, 5] with pretrained vision encoders [22, 16], using a vision projector to connect visual features to the LLMs.
The standard training process for MLLMs generally involves two stages: (1) a vision projector pretraining stage, where only the vision projector is trained to align visual features for the LLMs, and (2) an instruction tuning stage, where both the vision projector and the LLM are trained to perform visual instruction tasks. To reduce the extensive number of parameters in LLM tuning, low-rank adaptation (LoRA) [8] was proposed by injecting lightweight adapters into pretrained LLMs for efficient fine-tuning. The vision projector pretraining aligns visual and language feature spaces, while instruction tuning equips MLLMs to generate responses aligned with human instructions. The goal of these two stages is to equip MLLMs with sufficient visual cues and to ensure accurate instruction-based responses.
Despite the great success, existing MLLMs have two main limitations. On the one hand, in the vision projector pretraining stage, current methods primarily rely on high-level visual features, often overlooking low-level and fine-grained details, which limits visual comprehension. For example, most works[15, 2] use only the high-level semantic feature map for vision token projection, while some [24, 31] leverage multi-layer visual features yet still miss finer details. On the other hand, in the instruction-tuning stage, as downstream tasks grow in diversity and complexity, data conflicts in LoRA-based instruction tuning become more prominent  [4, 9]. To mitigate this, recent studies [4, 18] incorporate the Mixture of Experts (MoE) paradigm into LoRA modules, embedding multiple LoRA units within a single linear layer so that each unit can leverage its specific strengths. To better accommodate varied tasks, LoRA experts with different levels of granularity or customized routing strategies [18] have been introduced. However, this complex design not only increases implementation complexity but may also extend the time required for training and inference. We thus raise the question: Is there a simple and unified approach to flexibly represent any group of LoRA experts?
Figure 1: The two training stages of our framework. In the Vision Projector Pretraining stage, we incorporate the proposed Visual Cue Enhancement module (VCE) to enhance the final vision feature map by capturing detailed local information from multi-level visual feature maps. In the Visual Instruction Fine-Tuning stage, we use the proposed Dual Low-Rank Adaptation (Dual-LoRA) as the LLM adapter to learn task-specific knowledge while mitigating data conflicts, which is set trainable alone with VCE and vision projector.
In this paper, we address existing limitations by proposing Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA) to improve the vision projector pretraining stage and the instruction finetuning stage respectively. An overview of our overall training pipeline is shown in Figure 1. For the VCE module, we extend feature capture to multi-level vision feature maps instead of relying solely on high-level semantic feature maps. Specifically, for each patch in the high-level feature map, we extract neighboring patches from the middle-layer feature maps and combine them with the original patch to create a locally enhanced feature patch. This approach enhances local visual cues while maintaining computational efficiency. To address data conflicts during MLLM instruction tuning, we introduce Dual-LoRA, a unified adaptation module that decouples learning into a skill low-rank space and a task-activation low-rank space. The skill space captures specific knowledge in downstream tasks, while the task space activates prior knowledge based on specific instructions. These two spaces are then jointly mapped into a cohesive single space to generate features for the subsequent stages. This separation of skill and task spaces within a single LoRA module enables fine-grained control over skill knowledge activation, allowing for dense LoRA rank learning of skill knowledge while supporting sparse, task-specific responses.
Experiments on downstream tasks and general MLLM benchmarks demonstrate the effectiveness of our proposed method. Our contributions can be summarized as follows:
  • We introduce a lightweight Visual Cue Enhancement (VCE) module to enrich fine-grained visual features for vision projector pretraining.
  • We propose Dual Low-Rank Adaptation (Dual-LoRA), a simple yet efficient approach that mitigates data conflicts for effective Visual Instruction Fine-Tuning.
  • Experimental results on both downstream tasks and general MLLM benchmarks validate the effectiveness of our proposed method.
Figure 2: Illustration of our proposed components. The Visual Cue Enhancement (VCE) module uses the projected vision feature map $f^{\textbf{a}}$ selected during pretraining as the anchor feature, leveraging nearby patches in multi-layer feature maps $[f^{\textbf{ref}}]$ to enrich the local visual cues of $f^{\textbf{a}}$. The Dual Low-Rank Adaptation decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
2 Related Works
Multi-Modal Large Language Models (MLLMs) integrate pretrained vision encoders with pretrained large language models (LLMs), showcasing remarkable instruction-following and generalization capabilities. Early models like Flamingo and BLIP-2 successfully adapted LLMs for visual tasks, demonstrating strong zero-shot generalization and in-context learning abilities. Visual instruction tuning, a key technique in recent studies [14], has further enhanced MLLMs by using vision-language (VL) datasets and improving visual instruction-following data. Additionally, some research approaches vision as a language, employing textual annotations (e.g., location descriptions of bounding boxes) to train MLLMs effectively.
Enhancing Vision Feature Representation in MLLMs primarily involves vision encoder ensembling, resolution enhancement, and multi-level feature map fusion. For vision encoder ensembling, MMVP [25] utilizes a Mixture of Features (MoF) strategy to integrate image features from CLIP-ViT [22] and DINOv2 [20]. Similarly, MouSi [7] uses an ensemble technique to leverage the strengths of individual vision encoders, introducing a fusion network to unify outputs from different encoders, such as CLIP, DINOv2, and SAM. DeepSeekVL [19] adopts a hybrid vision encoder design, encoding images by combining SigLIP-L for low-resolution inputs and SAM-B for high-resolution inputs. LLaVA-UHD [29] organizes input images for efficient and scalable encoding, utilizing a compression module to further condense image tokens from visual encoders. Some approaches focus on fusing multi-level feature maps. For example, Cambrian-1 [24] introduces the Spatial Vision Aggregator, a dynamic, spatially-aware connector that integrates high-resolution visual features with LLMs while minimizing token count. While most of these methods require full fine-tuning of the LLM, we propose a lightweight visual cue enhancer for more efficient fine-tuning.
Low-Rank Adaptation (LoRA) [8] introduces a method that freezes pretrained model weights and injects trainable low-rank decomposition matrices into the linear layers of the Transformer architecture, significantly reducing the number of trainable parameters required for downstream tasks. Mixture of Experts in LoRA (MoE-LoRA) has gained popularity for addressing data conflicts in downstream task fine-tuning. MoE-LoRA typically employs three types of activation: sparse activation [4], where only the top-$k$ experts are activated; dense activation [27], where all experts are activated and re-weighted by a router; and rectified activation [10], where experts are dynamically activated and optimized to be sparse. For instance, LLaVA-MoLE selects only the top-1 expert for specific tasks. In dense activation, MoLE [27] treats each trained LoRA module at each layer as an individual expert, integrating learnable gating functions at each layer to control hierarchical weights. In rectified activation, RoDE [10] designs diverse experts and employs a rectified linear unit to optimize the router for sparse expert activation. These methods involve complex expert group configurations or routing strategies. In contrast, our approach offers a simple yet efficient adapter that effectively addresses data conflicts without requiring a complicated design.
Figure 3: Illustration of (a) VCE, (b) typical Mixture of Experts strategies for addressing data conflicts, and (c) Dual-LoRA. Mixture of Experts strategies include: Sparse activation [4], where only a subset of experts is activated; dense activation [27], where all experts contribute equally; and rectified activation [10], where all experts contribute with selective rectification.
3 Method
3.1 Preliminaries
Multi-Modal Large Language Models (MLLMs). A standard MLLM (e.g., LLaVA [14]) consists of an image feature encoder $V(\cdot)$, a vision feature projector $P(\cdot)$, and a large language model $G(\cdot)$. Given an image-question-answer (VQA) triplet $(I,T_{\text{que}},T_{\text{ans}})$, $V$ extracts image features, which are then projected into vision tokens $t_{\text{vision}}$ through $P$. These vision tokens are concatenated with question text tokens $t_{\text{que}}$, and the combined sequence is passed to $G$ to generate predicted answer tokens $t_{\text{pred}}$. Instruction tuning aims to optimize the parameters of $P$ and $G$ so that $t_{\text{pred}}$ closely matches the target answer tokens $t_{\text{ans}}$:
$\theta_{P},\theta_{G}$ $=\arg\min_{\theta_{P},\theta_{G}}\mathcal{L}(t_{\text{pred}},t_{\text{ans}}),$ (1)
$t_{\text{pred}}$ $=G(t_{\text{vision}},t_{\text{que}};\theta_{G}),$ (2)
$t_{\text{vision}}$ $=P(V(I);\theta_{P}).$ (3)
Here, $\theta_{P}$ and $\theta_{G}$ represent the trainable parameters of $P$ and $G$, respectively, and $\mathcal{L}(\cdot,\cdot)$ denotes the loss function, typically the cross-entropy loss.
Low-Rank Adaptation (LoRA). Due to the large number of parameters in $G$, optimizing $\theta_{G}$ can be highly resource-intensive. To address this, LoRA [8] has been proposed as a lightweight adapter injected into the linear layers of LLMs. LoRA significantly reduces the number of trainable parameters by decomposing the transformation of the input feature into a low-rank form. The LoRA module is represented by two parameter matrices, $A$ and $B$ with the same rank $r$. In the $i$-th injected layer, let $W_{i}$ denote the pre-trained weights and $x_{i}$ denote the input feature for this layer; then the output after applying the LoRA module can be expressed as:
$z_{i}$ $=W_{i}x+\frac{r}{\alpha}B_{i}A_{i}x,$ (4)
where $z_{i}$ is the output feature passed to the next layer, and $\alpha$ is a scaling factor. For convenience, we omit the indices $i$ in the following. With LoRA, the trainable parameters of $F$ are reduced to $\{\theta_{P},A,B\}$.
3.2 Framework Overview
We propose a two-stage approach for efficient Visual Instruction Tuning: (1) Enhanced Vision Projector Pretraining and (2) Dual Low-Rank Adaptation (Dual-LoRA) Visual Instruction Fine-Tuning, as illustrated in Figure 1. Specifically, to enable comprehensive vision information processing and effective execution of instruction-following downstream tasks, we propose two main components: Visual Cue Enhancement (VCE) to capture enriched visual cues by incorporating multi-level visual features, and Dual-LoRA to mitigate data conflicts by decoupling visual instruction tuning into skill-task dual low-rank space learning. The complete framework pipeline is illustrated in Figure 2, with further details provided in the following sections.
3.3 Multi-level Local Visual Cue Enhancement
Typical vision projectors primarily rely on high-level vision feature maps to align vision and language modalities. For example, LLaVA [14] and Qwen-VL [1] use the penultimate layer output of ViT, which captures high-level, language-aligned semantic features but may overlook finer, detail-oriented visual cues.
To address these limitations, we propose a lightweight Vision Cues Enhancement (VCE) module. This module uses the final vision feature map as an anchor feature and applies a single deformable cross-attention mechanism to neighboring patches across other layer feature maps for each patch of the anchor feature. An intuitive illustration is shown in Figure 3 (a).
Denote $F=[f]$ as the multi-level feature maps produced by the vision encoder $V$. We refer to the high-level feature map used for projection into vision tokens as the anchor feature $f^{\text{a}}$, while several intermediate layers are designated as reference feature maps $[f^{\text{ref}}]$. The vision encoder processes the input image $I$ through multiple blocks to extract these multi-level features $F$, formulated as:
[…]
where select-index specifies the index of reference feature maps.
To enhance the vision cues, we integrate local visual details from $[f^{\text{ref}}]$ into $f^{\text{a}}$. We utilize deformable attention [28] to efficiently compute cross-attention across multi-level feature maps, applied between $[f^{\text{ref}}]$ and $f^{\text{a}}$:
$f^{*}=\text{deform-attn}(f^{\text{a}},[f^{\text{ref}}];\theta_{\text{deform-attn}}),$ (6)
[…]
After obtaining the anchor feature $f^{\text{a}}$ and the enhanced visual cue $f^{*}$, we add them up and apply normalization to produce the final enhanced vision feature map $f_{\text{vision}}$. This feature map is subsequently projected into vision tokens $t_{\text{vision}}$ through the vision feature projector $P$:
$f_{\text{vision}}$ $=\text{norm}(f^{\text{a}}+f^{*}),$ (7)
$t_{\text{vision}}$ $=P(f_{\text{visual}};\theta_{P}).$ (8)
The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the multimodal model $G$ for further processing.
3.4 Dual Low Rank Adaptation
LoRA assumes that in large pre-trained models, the model parameters have a high degree of redundancy, with a low-rank structure that can be effectively represented through low-rank matrix approximation. However, as downstream tasks become more complex or diverse, data conflicts can arise due to the limited representational capacity of the low-rank structure [9, 4]. Some methods integrate Mixture of Experts into LoRA to better adapt to diverse downstream tasks. They propose expert modules with different granularities [10] and use various routing strategies [4, 27, 18, 10].
In contrast, we propose Dual Low-Rank Adaptation (Dual-LoRA), a unified approach that decouples adaptation into a task-skill dual low-rank structure, removing the need for complex expert splitting or intricate routing strategies. Figure 3 (b) and (c) illustrate the differences between our method and other strategies for mitigating data conflicts.
Proposition 1: For a set of LoRAs $\{B_{k}A_{k}\}_{k=1}^{K}$, each with rank 1, where $K$ is the total number of LoRAs, the representation space of a single LoRA with rank $K$ contains, and is at least as expressive as, the union of the representation spaces of the individual rank-1 LoRAs.
Corollary 1: For a set of LoRAs $\{B_{k}A_{k}\}_{k=1}^{K}$, the combined representation capability does not exceed that of a single LoRA with rank $\sum\text{rank}(\text{LoRA}_{k})$.
Corollary 2: For a LoRA $BA$ with rank $r$, the LoRA $B(A\cdot\sigma(T))$ can be decomposed into any set of LoRA groups $\{B_{k}A_{k}\}_{k=1}^{K}$, provided that the constraint $r\geq\sum_{k=1}^{K}\text{rank}(B_{k}A_{k})$ holds. Here, $T$ is a matrix with the same shape as $A$, and $\sigma$ is a non-linear activation function.
The proofs of Proposition 1 and its corollary are provided in Supplementary Materials A. Inspired by these findings, we aim to explore the feasibility of using a single larger-rank LoRA to achieve the same effect mixture of LoRA experts.
3.4.1 Skill Low-Rank Space
We propose the construction of a skill low-rank space to encapsulate specific knowledge for downstream tasks while ensuring domain knowledge consistency. According to Corollary 1, a single LoRA with equivalent parameters can theoretically represent any configuration of LoRA experts across various ranks without loss of expressive power. Let $\mathcal{S}$ represent the skill space of the unified LoRA, modeled by a parameter metric $S$.
In practice, however, a split LoRA expert group often outperforms a single, larger LoRA [10]. This can be attributed to variations in routing strategies and initialization across individual LoRAs. For example, LLaVA-MoLE [4] activates the top-k experts for targeted learning, while RoDE [10] employs heterogeneous LoRA experts, each initialized with distinct distributions. The complex training process, particularly the routing score, leads to divergence in the representational spaces learned by different LoRAs over time. The complex training process, especially the routing score, leads to divergence in the representational spaces learned by different LoRAs over time. Intuitively, even for the same input, each expert may experience unique gradients due to its intrinsic numerical differences and routing scores.
This observation suggests the need for an additional space to mediate between the instruction task and the skill space $\mathcal{S}$, allowing $S$ to be allocated with numerical differences according to instruction variance.
3.4.2 Rectified Skill Low-Rank Space
Based on the skill low-rank space, we aim to activate prior knowledge in the task space guided by specific instructions. According to Corollary 2, additional space can be used to map the skill space $\mathcal{S}$, enabling the simulation of sparsely activated experts that respond to varying instructional tasks. We use the parameter matrix $T$ to modulate $S$, which is mapped through a non-linear activation $\sigma$ to achieve the effect of rank rectification. In this paper, we use the ReLU [11] activation function due to its sparsity-inducing properties and ease of optimization. The rectified, decomposed feature space, mapped from the skill and task spaces, can then be expressed as:
$D(x)$ $=\frac{r}{\alpha}B(Sx\cdot\sigma(Tx)),$ (9)
where $D$ denotes the Dual-LoRA module. To smooth the distribution of the skill space and improve stability in subsequent computations, we normalize the skill space using layer normalization, which can be easily implemented thanks to the unified single LoRA design. With normalization applied to the skill space, Equation 9 can be revised as:
$D(x)$ $=\frac{r}{\alpha}B(\text{Norm}(Sx)\cdot\sigma(Tx)).$ (10)
Finally, we combine the adaptation feature generated by $D$ with the feature generated from the pre-trained weight $W$ to obtain the output feature $z$:
$z$ $=W(x)+\frac{r}{\alpha}B(\text{Norm}(Sx)\cdot\sigma(Tx)).$ (11)
3.5 Training
Our method consists of two training stages, as shown in Figure 1. In the vision projector pretraining stage, we train the parameters of the vision projection modules, i.e., $\{\theta_{P},\theta_{\text{deform-attn}}\}$, using descriptive image-caption pairs. In the subsequent visual instruction fine-tuning stage, we enable additional training for the Dual-LoRA parameters, where $\{\theta_{P},\theta_{\text{deform-attn}},S,T,B\}$ are set to be trainable, leveraging the full range of downstream tasks.
Method Ingredient Recognition Recipe Generation Nutrition Estimation (pMAE $\downarrow$ )
IoU $\uparrow$ F1 $\uparrow$ SacreBLEU $\uparrow$ Rouge-L $\uparrow$ mass cal fat protein carb avg
vanilla LoRA [8] 23.2 34.1 12.4 40.1 46.2 45.5 57.1 53.4 48.7 50.2
MoE-LoRA (top-2) [27] 22.9 33.8 12.7 40.2 45.56 45.8 56.9 54.4 48.0 50.1
MoE-LoRA (softmax) [27] 22.7 33.5 12.5 40.0 45.3 45.5 58.1 53.7 47.5 50.0
RoDE [10] 23.6 34.6 13.8 41.4 45.8 47.6 58.5 54.4 50.4 51.3
Dual-LoRA 24.2 35.2 14.8 42.1 46.1 46.2 56.8 52.2 48.7 49.9
Dual-LoRA + VCE 24.5 35.5 14.7 42.2 44.9 45.0 56.8 51.2 47.5 49.1
Table 1: Performance comparison with state-of-the-art Mixture of Experts in LoRA approaches on UniFood dataset. For fair comparison, all methods use the same base pretrained MLLM, LLaVA-1.5-7B. The best performance is highlighted in bold, and the second-best is underlined.
[…]
InstructBLIP [6] Vicuna-7B EVA-CLIP ViT-G 224 36.0 60.9 58.8 - 26.2
IDEFICS [3] LLaMA-7B OpenCLIP ViT-H/14 224 48.2 - 44.5 - -
Shikra [3] Vicuna-7B CLIP ViT-L/14 224 58.8 - - - -
Qwen-VL [1] Qwen-7B OpenCLIP ViT-bigG 448 38.2 - 62.3 - -
Qwen-VL-Chat [1] Qwen-7B OpenCLIP ViT-bigG 448 60.6 65.4 - - -
LLaVA-1.5 [15] Vicuna-7B CLIP ViT-L/14 336 64.3 - 65.4 85.9 31.1
LLaVA-1.5-LoRA [15] Vicuna-7B CLIP ViT-L/14 336 66.1 - 67.9 86.4 30.2
Dual-LoRA Vicuna-7B CLIP ViT-L/14 336 65.1 64.5 67.4 86.8 31.5
Dual-LoRA+VCE Vicuna-7B CLIP ViT-L/14 336 65.6 64.7 68.1 87.2 32.1
Table 2: Performance comparison with state-of-the-art MLLMs on general benchmarks. The best performance is highlighted in bold, and the second-best is underlined.
4 Experiments
4.1 Model and Optimizer Configuration
For the base pre-trained multimodal large language model, we use the widely adopted open-source model LLaVA-1.5-7B [14]. LLaVA-1.5-7B integrates a CLIP ViT-L [22] with an image resolution of 336px and a patch size of 14, a two-layer MLP projector to map visual features into tokens, and Vicuna v1.5 [5] as the language model. In our setup, we omit any special tokens, such as image indicator tokens, to enclose visual tokens. For the VCE module configuration, we use the outputs from the 2nd-to-last, 8th-to-last, 14th-to-last, and 20th-to-last layers of CLIP ViT-L as reference feature maps (see Section 3.3 for details). In the Dual-LoRA setting, we set the LoRA dropout rate to 0.05.
Our VCE module requires only 5.52 MB of memory, compared to the 40.02 MB needed for the vision projector in LLaVA-1.5-7B. Furthermore, our Dual-LoRA task consumes approximately 1.5 MB per LoRA rank when injected solely into the query projection and value projection layers. For optimization, we use the Adam optimizer across all trainable parameters. During the pretraining stage, we set a global batch size of 256 and a learning rate of 0.001 with no weight decay. In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LMM adapter and 0.00002 for the vision feature projector, respectively.
4.2 Evaluation on Downstream Tasks
4.2.1 Experiment Setup
Dataset and Downstream Tasks. We begin by evaluating our proposed method on downstream fine-tuning tasks using the UniFood dataset [10], a multi-task dataset with 100,000 samples in the food domain. The tasks in UniFood include ingredient recognition, recipe generation, and nutrition estimation. For ingredient recognition, we use Intersection over Union (IoU) and F1-score as evaluation metrics. For recipe generation, we assess performance with SacreBLEU [21] and Rouge-L [13]. For nutrition estimation, we use the percentage of mean average error (pMAE) [32] as the metric.
Baselines. We select several Mixture of Experts (MoE) methods within the LoRA framework: a sparse activation method, MoE-LoRA (top-2), which activates the top-2 experts; a dense activation method, MoE-LoRA (softmax), which uses a softmax router to engage all experts; and a rectified activation method, RoDE [10], with rectified diverse expert activation. Both MoE-LoRA (top-2) and MoE-LoRA (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$. Additionally, we include a standard LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
Implementation Details. We conduct our experiments on a setup with 8 RTX 4090 GPUs. The LLM adapters are integrated only within the linear layers of the query and value projectors in the transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.
4.2.2 Performance Comparison
Table 1 presents the results of our method compared to baseline methods on the UniFood dataset. As shown, our Dual-LoRA method consistently outperforms the sparse expert activation method (Top-2), the dense expert activation method (Softmax), and the rectified expert activation method (RoDE) across all tasks. Additionally, incorporating the proposed VCE module with Dual-LoRA (i.e., Dual-LoRA+VCE) manages to further boost performance for both ingredient recognition and nutrition estimation tasks. These results demonstrate the effectiveness of our proposed Dual-LoRA and VCE components for instruction fine-tuning in downstream tasks.
Figure 4: Qualitative results. Dual-LoRA + VCE produces more accurate and knowledge-consistent answers, whereas the LLaVA-1.5-7b-LoRA generates knowledge-inconsistent responses.
Figure 5: Performance vs. trainable LLM adapter parameters for vanilla LoRA and our proposed Dual-LoRA on the UniFood dataset. The adapters are integrated into the query and value projectors of the transformer blocks. To eliminate the influence of trainable parameters beyond the LLM adapter, only the LLM adapter is set as trainable.
4.3 Evaluation on General MLLM Benchmarks
4.3.1 Experiment Setup
MLLM Benchmarks. For the general benchmark evaluation, following LLaVA [14], we use a 558K subset of the LAION-CC-SBU dataset [23] and the llava-v1.5-mix665k dataset [14] for instruction fine-tuning. Four widely used MLLM benchmarks are used for evaluation, including MMBench [17], SEED-Bench [12], and LLaVA-Bench In-the-${}^{\text{W}}$ ) [14]. The first three benchmarks assess various MLLM capabilities, such as perceptual understanding and visual reasoning, through binary yes/no questions (MME) or multiple-choice questions (MMBench and SEED-Bench). We use the image-only subset of SEED-Bench (SEEDI), a popular choice for many image-based MLLMs [2, 30].
Implementation Details. Our experiments are conducted on a setup with 4 $\times$ A100 GPUs (80GB). The selected layers for LLM adapter injection match those used in LLaVA-1.5-7B-LoRA [14], with the adapter rank $r$ set to 128 and scaling factor $\alpha$ set to 256 in Equation 11.
4.3.2 Performance Comparison
The results of our proposed method, alongside state-of-the-art MLLMs, are presented in Table 2. The Dual-LoRA + VCE model consistently ranks among the top two across all benchmark tests. Compared to other methods, this model demonstrates robust performance across multiple benchmarks, indicating excellent generalization and strong adaptability to a variety of tasks. These findings validate the effectiveness of the VCE module and Dual-LoRA, enabling the model to achieve superior results on general visual understanding tasks.
VCE Dual-LoRA Ingredient Recognition Recipe Generation
IoU F1 SacreBLEU Rouge-L
23.2 34.1 12.4 40.1
24.2 35.2 14.8 42.1
23.3 34.2 12.6 40.5
24.5 35.5 14.7 42.2
Table 3: Ablation study of the main components of our proposed method: Visual Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). The baseline model vanilla LoRA with LLaVA-1.5-7B is presented in the first row.
Skill space @Norm Task space @ Non-linear Ingredient Recognition Recipe Generation
IoU F1 SacreBLEU Rouge-L
21.4 32.0 13.1 41.1
22.6 33.2 14.1 41.2
21.9 32.6 13.4 41.3
23.0 33.7 14.4 41.7
24.2 35.2 14.8 42.1
Table 4: Ablation studies for the design of Dual-LoRA. “@Norm” indicates the use of layer normalization to process the output $Sx$, while “@Non-linear” denotes the use of a non-linear activation function to process the output $Tx$. The color $\blacksquare$ represents a frozen vision projector, whereas $\blacksquare$ indicates it is trainable.
Rank LMM Adapter LMM Adapter Memory Size Ingredient Recognition Recipe Generation
IoU F1 SacreBLEU Rouge-L
var1 32 LoRA 32M 19.2 29.7 10.4 41.1
var2 Dual-LoRA 48M 22.7 +18.1% 33.6 +13.3% 12.9 +24.1% 41.3 +0.3%
var3 64 LoRA 64M 20.6 31.3 12.7 41.3
var4 Dual-LoRA 96M 23.0 +11.7% 33.7 +7.7% 14.4 +13.1% 41.7 +0.9%
var5 128 LoRA 128M 21.6 32.5 13.0 41.3
var6 Dual-LoRA 192M 23.6 +8.9% 34.4 +5.9% 14.8 +14.3% 42.9 +3.6%
Table 5: Performance and memory usage of the LoRA adapter under varying rank settings on the UniFood dataset. To prevent the effect of vision projector, only the LoRA adapter is set to be trainable.
4.4 Qualitative Results
Figure 4 presents a comparison between our proposed method and the vanilla LoRA fine-tuned version of LLaVA-1.5-7B on food-related downstream tasks. The vanilla LoRA demonstrates knowledge inconsistency, with the ingredient list failing to align with the generated cooking instructions. In contrast, our method produces more accurate and knowledge-consistent answers, ensuring the ingredient list aligns with the cooking instructions. These results visually underscore the effectiveness of our approach.
4.5 Ablation Study
In this section, we conduct ablation studies to verify different components of our method on UniFood dataset. The setup is the same as Section 4.2.1.
VCE and Dual-LoRA. We conducted an ablation study to analyze the impact of different projector and adapter configurations on the performance of our model in ingredient recognition and recipe generation. The results are shown in Table 3. This ablation study shows that both multi-level local visual feature enhancement (VCE) and Dual-LoRA positively impact performance. When both are enabled, the best results are achieved for ingredient recognition and recipe generation tasks, indicating that the combination of these methods is more effective than using either one alone.
Dual-LoRA design. We further investigate the design choices of our Dual-LoRA. Table 4 presents the results of an ablation study on the normalization of $S$ and the non-linear activation of $T$ (see Section 3.4 for details). This ablation study shows that removing the non-linear activation in the task space or the normalization in the skill space leads to a performance decline. When both are removed, the performance of Dual-LoRA reaches its lowest level.
4.6 Efficiency Analysis
As listed in Table 5, we present the results of the vanilla LoRA and Dual-LoRA across three configurations of total rank values (32, 64, and 128). The results indicate that Dual-LoRA consistently outperforms the vanilla LoRA counterparts across all rank settings (32, 64, and 128) with only a slightly higher number of parameters. Notably, Dual-LoRA achieves superior performance even with fewer trainable parameters than LoRA with higher ranks, such as Var2-Dual-LoRA vs. Var3-LoRA and Var5-LoRA, as well as Var4-Dual-LoRA vs. Var5-LoRA, which is also visualized in Figure 5. We also observe that Dual-LoRA shows relatively larger improvements over vanilla LoRA when the rank size is small. We attribute this to data conflicts being more obvious with fewer trainable parameters, and the incorporation of task space in Dual-LoRA helps mitigate this issue. These findings highlight the effectiveness of Dual-LoRA in addressing data conflicts when training multiple tasks in downstream applications.
5 Conclusion
In this paper, we explore efficient visual instruction fine-tuning for multi-modal large language models (MLLMs) by leveraging pre-trained models and utilizing only a few adapter parameters. Visual instruction fine-tuning of MLLMs faces two key challenges: insufficient capture of visual cues and conflicts among various instruction tasks due to the limited adapter parameters. To address these challenges, we propose a framework comprising Vision Cues Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). For Vision cue enhancement, we use multi-level feature maps to enrich the extracted vision cues. For Dual Low-Rank Adaptation, we introduce a rectified skill-space mapping technique, enabling the model to learn complex knowledge and respond specifically to instructions. Dual Low-Rank Adaptation possesses the expressive power of any ensemble of experts with a combined rank lower than its own. Evaluations on both general visual instruction benchmarks and domain-specific benchmarks demonstrate the effectiveness of our proposed method.

### v2

@@
 Abstract
-Fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity.
+Parameter-efficient fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity.
 To address these issues, we propose an efficient fine-tuning framework with two novel approaches: Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA).
@@
 To mitigate this, recent studies [4, 18] incorporate the Mixture of Experts (MoE) paradigm into LoRA modules, embedding multiple LoRA units within a single linear layer so that each unit can leverage its specific strengths.
-To better accommodate varied tasks, LoRA experts with different levels of granularity or customized routing strategies [18] have been introduced.
+To better accommodate varied tasks, LoRA experts with different levels of granularity [10] or customized routing strategies [18] have been introduced.
 However, this complex design not only increases implementation complexity but may also extend the time required for training and inference.
@@
 This separation of skill and task spaces within a single LoRA module enables fine-grained control over skill knowledge activation, allowing for dense LoRA rank learning of skill knowledge while supporting sparse, task-specific responses.
-Experiments on downstream tasks and general MLLM benchmarks demonstrate the effectiveness of our proposed method. Our contributions can be summarized as follows:
+Experiments conducted on downstream tasks within the food domain, as well as on several general MLLM benchmarks, highlight the effectiveness of our proposed method.
+Our contributions can be summarized as follows:
   • We introduce a lightweight Visual Cue Enhancement (VCE) module to enrich fine-grained visual features for vision projector pretraining.
@@
 The Visual Cue Enhancement (VCE) module uses the projected vision feature map $f^{\textbf{a}}$ selected during pretraining as the anchor feature, leveraging nearby patches in multi-layer feature maps $[f^{\textbf{ref}}]$ to enrich the local visual cues of $f^{\textbf{a}}$.
-The Dual Low-Rank Adaptation decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
+The Dual Low-Rank Adaptation (Dual-LoRA) decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
 2 Related Works
@@
 $t_{\text{vision}}$ $=P(f_{\text{visual}};\theta_{P}).$ (8)
-The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the multimodal model $G$ for further processing.
+The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the large language modal model $G$ for further processing.
+The structure of the adaptive $G$ is detailed in the next section.
 3.4 Dual Low Rank Adaptation
@@
 Corollary 1: For a set of LoRAs $\{B_{k}A_{k}\}_{k=1}^{K}$, the combined representation capability does not exceed that of a single LoRA with rank $\sum\text{rank}(\text{LoRA}_{k})$.
-Corollary 2: For a LoRA $BA$ with rank $r$, the LoRA $B(A\cdot\sigma(T))$ can be decomposed into any set of LoRA groups $\{B_{k}A_{k}\}_{k=1}^{K}$, provided that the constraint $r\geq\sum_{k=1}^{K}\text{rank}(B_{k}A_{k})$ holds.
+Corollary 2: For a LoRA $BA$ with rank $r$, the LoRA $B(A\odot\sigma(T))$ can be decomposed into any set of LoRA groups $\{B_{k}A_{k}\}_{k=1}^{K}$, provided that the constraint $r\geq\sum_{k=1}^{K}\text{rank}(B_{k}A_{k})$ holds.
 Here, $T$ is a matrix with the same shape as $A$, and $\sigma$ is a non-linear activation function.
@@
 The complex training process, particularly the routing score, leads to divergence in the representational spaces learned by different LoRAs over time.
-The complex training process, especially the routing score, leads to divergence in the representational spaces learned by different LoRAs over time.
 Intuitively, even for the same input, each expert may experience unique gradients due to its intrinsic numerical differences and routing scores.
@@
 The rectified, decomposed feature space, mapped from the skill and task spaces, can then be expressed as:
-$D(x)$ $=\frac{r}{\alpha}B(Sx\cdot\sigma(Tx)),$ (9)
+$D(x)$ $=\frac{r}{\alpha}B(Sx\odot\sigma(Tx)),$ (9)
 where $D$ denotes the Dual-LoRA module.
@@
 With normalization applied to the skill space, Equation 9 can be revised as:
-$D(x)$ $=\frac{r}{\alpha}B(\text{Norm}(Sx)\cdot\sigma(Tx)).$ (10)
+$D(x)$ $=\frac{r}{\alpha}B(\text{Norm}(Sx)\odot\sigma(Tx)).$ (10)
 Finally, we combine the adaptation feature generated by $D$ with the feature generated from the pre-trained weight $W$ to obtain the output feature $z$:
-$z$ $=W(x)+\frac{r}{\alpha}B(\text{Norm}(Sx)\cdot\sigma(Tx)).$ (11)
+$z$ $=Wx+\frac{r}{\alpha}B(\text{Norm}(Sx)\odot\sigma(Tx)).$ (11)
 3.5 Training
@@
 During the pretraining stage, we set a global batch size of 256 and a learning rate of 0.001 with no weight decay.
-In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LMM adapter and 0.00002 for the vision feature projector, respectively.
+In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LLM adapter and 0.00002 for the vision feature projector, respectively.
 4.2 Evaluation on Downstream Tasks
@@
 Both MoE-LoRA (top-2) and MoE-LoRA (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$.
-Additionally, we include a standard LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
+Additionally, we include a vanilla LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
 Implementation Details.
 We conduct our experiments on a setup with 8 RTX 4090 GPUs.
-The LLM adapters are integrated only within the linear layers of the query and value projectors in the transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.
+The LLM adapters are integrated only within the linear layers of the query and value projectors in the LLM transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.
 4.2.2 Performance Comparison

### v3

@@
-Visual Cue Enhancement and Dual Low-Rank Adaptation for Efficient Visual Instruction Fine-Tuning
+From Holistic to Localized: Local Enhanced Adapters for Efficient Visual Instruction Fine-Tuning
 Abstract
-Parameter-efficient fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity.
-To address these issues, we propose an efficient fine-tuning framework with two novel approaches: Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA).
-VCE enhances the vision projector by integrating multi-level visual cues, improving the model’s ability to capture fine-grained visual features.
-Dual-LoRA introduces a dual low-rank structure for instruction tuning, decoupling learning into skill and task spaces to enable precise control and efficient adaptation across diverse tasks.
-Our method simplifies implementation, enhances visual comprehension, and improves adaptability.
-Experiments on both downstream tasks and general benchmarks demonstrate the effectiveness of our proposed approach.
+Efficient Visual Instruction Fine-Tuning (EVIT) seeks to adapt Multimodal Large Language Models (MLLMs) to downstream tasks with minimal computational overhead.
+However, as task diversity and complexity increase, EVIT faces significant challenges in resolving data conflicts.
+To address this limitation, we propose the Dual Low-Rank Adaptation (Dual-LoRA), a holistic-to-local framework that enhances the adapter’s capacity to address data conflict through dual structural optimization.
+Specifically, we utilize two subspaces: a skill space for stable, holistic knowledge retention, and a rank-rectified task space that locally activates the holistic knowledge.
+Additionally, we introduce Visual Cue Enhancement (VCE), a multi-level local feature aggregation module designed to enrich the vision-language projection with local details.
+Our approach is both memory- and time-efficient, requiring only 1.16$\times$ the inference time of the standard LoRA method (with injection into the query and value projection layers), and just 73% of the inference time of a 4-expert LoRA-MoE.
+Extensive experiments on various downstream tasks and general MLLM benchmarks validate the effectiveness of our proposed methods.
+Our project page are publicly available at https://github.com/pengkun-jiao/Dual-LoRA
+Figure 1: The motivation behind the proposed approach.
+(a) Humans perceive visual signals from both holistic and localized perspectives, considering the overall image structure and local details.
+(b) Humans acquire holistic knowledge, with specific parts (localized) applied to particular tasks.
+For example, one might first acquire holistic knowledge of cooking and then apply specific knowledge, e.g.
+ingredient recognition, to source ingredients.
 1 Introduction
-Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across a wide range of tasks, showcasing versatility and effectiveness in diverse applications.
-Existing MLLMs [14, 1, 15] typically integrate pretrained Large Language Models (LLMs) [26, 5] with pretrained vision encoders [22, 16], using a vision projector to connect visual features to the LLMs.
-The standard training process for MLLMs generally involves two stages: (1) a vision projector pretraining stage, where only the vision projector is trained to align visual features for the LLMs, and (2) an instruction tuning stage, where both the vision projector and the LLM are trained to perform visual instruction tasks.
-To reduce the extensive number of parameters in LLM tuning, low-rank adaptation (LoRA) [8] was proposed by injecting lightweight adapters into pretrained LLMs for efficient fine-tuning.
-The vision projector pretraining aligns visual and language feature spaces, while instruction tuning equips MLLMs to generate responses aligned with human instructions.
-The goal of these two stages is to equip MLLMs with sufficient visual cues and to ensure accurate instruction-based responses.
-Despite the great success, existing MLLMs have two main limitations.
-On the one hand, in the vision projector pretraining stage, current methods primarily rely on high-level visual features, often overlooking low-level and fine-grained details, which limits visual comprehension.
-For example, most works[15, 2] use only the high-level semantic feature map for vision token projection, while some [24, 31] leverage multi-layer visual features yet still miss finer details.
-On the other hand, in the instruction-tuning stage, as downstream tasks grow in diversity and complexity, data conflicts in LoRA-based instruction tuning become more prominent  [4, 9].
-To mitigate this, recent studies [4, 18] incorporate the Mixture of Experts (MoE) paradigm into LoRA modules, embedding multiple LoRA units within a single linear layer so that each unit can leverage its specific strengths.
-To better accommodate varied tasks, LoRA experts with different levels of granularity [10] or customized routing strategies [18] have been introduced.
-However, this complex design not only increases implementation complexity but may also extend the time required for training and inference.
-We thus raise the question: Is there a simple and unified approach to flexibly represent any group of LoRA experts?
-Figure 1: The two training stages of our framework.
-In the Vision Projector Pretraining stage, we incorporate the proposed Visual Cue Enhancement module (VCE) to enhance the final vision feature map by capturing detailed local information from multi-level visual feature maps.
-In the Visual Instruction Fine-Tuning stage, we use the proposed Dual Low-Rank Adaptation (Dual-LoRA) as the LLM adapter to learn task-specific knowledge while mitigating data conflicts, which is set trainable alone with VCE and vision projector.
-In this paper, we address existing limitations by proposing Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA) to improve the vision projector pretraining stage and the instruction finetuning stage respectively.
-An overview of our overall training pipeline is shown in Figure 1.
-For the VCE module, we extend feature capture to multi-level vision feature maps instead of relying solely on high-level semantic feature maps.
-Specifically, for each patch in the high-level feature map, we extract neighboring patches from the middle-layer feature maps and combine them with the original patch to create a locally enhanced feature patch.
-This approach enhances local visual cues while maintaining computational efficiency.
-To address data conflicts during MLLM instruction tuning, we introduce Dual-LoRA, a unified adaptation module that decouples learning into a skill low-rank space and a task-activation low-rank space.
-The skill space captures specific knowledge in downstream tasks, while the task space activates prior knowledge based on specific instructions.
-These two spaces are then jointly mapped into a cohesive single space to generate features for the subsequent stages.
-This separation of skill and task spaces within a single LoRA module enables fine-grained control over skill knowledge activation, allowing for dense LoRA rank learning of skill knowledge while supporting sparse, task-specific responses.
-Experiments conducted on downstream tasks within the food domain, as well as on several general MLLM benchmarks, highlight the effectiveness of our proposed method.
-Our contributions can be summarized as follows:
-  • We introduce a lightweight Visual Cue Enhancement (VCE) module to enrich fine-grained visual features for vision projector pretraining.
-  • We propose Dual Low-Rank Adaptation (Dual-LoRA), a simple yet efficient approach that mitigates data conflicts for effective Visual Instruction Fine-Tuning.
-  • Experimental results on both downstream tasks and general MLLM benchmarks validate the effectiveness of our proposed method.
-Figure 2: Illustration of our proposed components.
-The Visual Cue Enhancement (VCE) module uses the projected vision feature map $f^{\textbf{a}}$ selected during pretraining as the anchor feature, leveraging nearby patches in multi-layer feature maps $[f^{\textbf{ref}}]$ to enrich the local visual cues of $f^{\textbf{a}}$.
-The Dual Low-Rank Adaptation (Dual-LoRA) decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
+Multimodal Large Language Models (MLLMs) have demonstrated impressive performances across a wide range of tasks, showcasing versatility and effectiveness in diverse applications [16, 15, 14].
+Existing MLLMs [20, 3, 21] typically integrate a pretrained Large Language Models (LLM) [32, 6] with a pretrained vision encoder [28, 22], using a vision projector to connect the vision encoder and LLM.
+Vision Instruction Tuning [20] enables the practical application of MLLMs in downstream tasks.
+However, full fine-tuning MLLMs can be highly resource-intensive.
+Efficient Vision Instruction Tuning (EVIT) addresses this challenge by tuning only the vision feature projector while integrating a small number of additional trainable parameters (adapters) into the LLM [8].
+For example, Low-Rank Adaptation (LoRA) [11] leverages low-rank approximations to replace full fine-tuning of LMM, significantly reducing computational overhead.
+Figure 2: Comparison between the proposed Dual-LoRA and existing methods.
+(a) Mainstream MLLMs, e.g.
+LLaVA [20] and Qwen-VL [3], project the high-level visual feature map.
+(b) Our Visual Cue Enhancement module enhances high-level features by aggregating local information from multi-level feature maps.
+LoRA-MoE methods mitigate data conflicts by enabling localized responses, i.e., experts activation, tailored to several activation strategies: (c) Sparse Activation[5], where only the top-k experts are activated; (d) Dense Activation[33], where all experts are activated; and Rectified Activation [13], where multiple heterogeneous experts are dynamically activated.
+In contrast, our (f) Dual Low-Rank Adaptation rectifies a holistic knowledge (skill) space with an additional task space, which is fully differentiable, capable of learning any local response, and more structurally efficient.
+However, existing EVIT methods face a critical limitation: as downstream tasks grow more diverse and complex, data conflicts in LoRA-based instruction tuning become increasingly pronounced [5, 12], ultimately impairing performance.
+An intriguing observation is that when tuning for a variety of food-related tasks, LoRA fine-tuning can result in knowledge-inconsistent answers.
+For example, some ingredients may be misaligned between food ingredient recognition and recipe generation, as illustrated in Fig.
+4.
+To address this, recent studies suggest integrating LoRA within a Mixture of Experts (LoRA-MoE) paradigm [5, 23, 33].
+LoRA-MoE adapts to diverse tasks by activating specific experts, which can be viewed as a process of localized adaptation, effectively mitigating data conflicts.
+However, the LoRA-MoE method requires a complex design to balance activation strategies, trainable parameters, and task complexity [13, 12].
+Moreover, it incurs significantly higher time consumption compared to vanilla LoRA when multiple LoRA experts are activated.
+To address the high time consumption of experts polling in LoRA-MoE, we propose a structurally efficient dual-space adaptation design that transitions from holistic domain knowledge learning to localized, task-specific adaptation.
+Additionally, we emphasize the role of localized adaptation in visual perception.
+A high-level overview of our motivation is illustrated in Figure 1.
+This design embodies a cognitively inspired learning paradigm: it first consolidates generalized skill representations and then dynamically routes task-critical knowledge components during inference, mirroring the transition in human cognition from holistic to localized processing.
+Specifically, we propose Dual Low-Rank Adaptation (Dual-LoRA), which strategically decouples adaptation learning into: 1) skill space learning for holistic domain knowledge acquisition, and 2) task space learning for instruction-conditioned task adaptation.
+The skill space is parameterized using a low-rank matrix optimized stably for cross-task knowledge consolidation, while the task space employs a rank-rectified matrix that dynamically modulates the skill space for localized task-specific adaptation.
+To enhance local details for vision feature projection, we further propose Visual Cue Enhancement (VCE), which enhances high-level visual features with local information aggregated from multiple feature maps.
+VCE remains computationally efficient by focusing on a limited number of local neighboring regions within each feature map.
+Figure 2 illustrates the structurally efficient design of our approach.
+Our VCE module is only 5.53 MB in size, and Dual-LoRA outperforms LoRA-MoE methods while using the same number of trainable parameters.
+The full approach requires just 1.16$\times$ the inference time of vanilla LoRA, and only 73% of the inference time of 4-experts LoRA-MoE methods.
+Extensive experiments across a range of downstream tasks—including UniFood, ScienceQA, Flickr30k, and general MLLM benchmarks—demonstrate the effectiveness of our proposed methods.
+Our main contributions can be summarized as follows:
+  • We design a time- and memory-efficient holistic-to-localized approach for EVIT. Extensive experiments demonstrate the effectiveness of our proposed methods.
+  • We introduce a lightweight Visual Cue Enhancement (VCE) module to enrich fine-grained visual features for efficient visual instruction fine-tuning.
+  • We propose Dual Low-Rank Adaptation (Dual-LoRA), an efficient approach that mitigates data conflicts for instruction fine-tuning.
 2 Related Works
-Multi-Modal Large Language Models (MLLMs) integrate pretrained vision encoders with pretrained large language models (LLMs), showcasing remarkable instruction-following and generalization capabilities.
-Early models like Flamingo and BLIP-2 successfully adapted LLMs for visual tasks, demonstrating strong zero-shot generalization and in-context learning abilities.
-Visual instruction tuning, a key technique in recent studies [14], has further enhanced MLLMs by using vision-language (VL) datasets and improving visual instruction-following data.
-Additionally, some research approaches vision as a language, employing textual annotations (e.g., location descriptions of bounding boxes) to train MLLMs effectively.
+Multimodal Large Language Models (MLLMs) harness the powerful reasoning capabilities of Large Language Models (LLMs) [32], driving significant advancements in vision tasks.
+Early works, such as Flamingo [1], resample visual features and integrate them into attention-based adapter layers within the LLM.
+BLIP-2 [18] introduces a Q-Former, which simultaneously handles cross-modal representation learning and generative learning.
+More recently, Vision-Language Instruction Tuning approaches, e.g.
+Instruct-BLIP [7], LLaVA [20], and Mini-GPT4 [37], have curated large-scale, high-quality multimodal instruction data to enhance MLLMs’ ability to follow instructions.
+However, fully tuning MLLMs is highly resource-intensive due to the vast number of parameters in LLMs, which hinders the application of MLLMs to downstream tasks.
+Low-Rank Adaptation (LoRA)[11] adapts the model by introducing low-rank matrices while keeping the original model parameters unchanged.
+Specifically, LoRA adds low-rank matrices to certain layers of the model (e.g., attention layers, feed-forward layers, etc.), enabling the model to fine-tune more efficiently for specific tasks.
+Mixture of Experts in LoRA (LoRA-MoE)[5, 13, 33] has gained popularity for addressing data conflicts during downstream task fine-tuning.
+LoRA-MoE methods can be grouped by the expert activation techniques: sparse activation [5], where only the top-k experts are activated; dense activation [33], where all experts are activated and re-weighted by a router; and rectified activation [13], where experts are dynamically activated and optimized for sparsity.
+RoDE [13] further introduces heterogeneous experts, where different sets of parameters are used for tasks of varying complexity.
+Figure 2 compares LoRA-MoE methods with our approach.
+LoRA-MoE activates multiple LoRA adapters, increasing inference time, whereas our method employs a single Dual-LoRA for greater inference efficiency.
 Enhancing Vision Feature Representation in MLLMs primarily involves vision encoder ensembling, resolution enhancement, and multi-level feature map fusion.
@@
 For example, Cambrian-1 [30] introduces the Spatial Vision Aggregator, a dynamic, spatially-aware connector that integrates high-resolution visual features with LLMs while minimizing token count.
-While most of these methods require full fine-tuning of the LLM, we propose a lightweight visual cue enhancer for more efficient fine-tuning.
-Low-Rank Adaptation (LoRA) [8] introduces a method that freezes pretrained model weights and injects trainable low-rank decomposition matrices into the linear layers of the Transformer architecture, significantly reducing the number of trainable parameters required for downstream tasks.
-Mixture of Experts in LoRA (MoE-LoRA) has gained popularity for addressing data conflicts in downstream task fine-tuning.
-MoE-LoRA typically employs three types of activation: sparse activation [4], where only the top-$k$ experts are activated; dense activation [27], where all experts are activated and re-weighted by a router; and rectified activation [10], where experts are dynamically activated and optimized to be sparse.
-For instance, LLaVA-MoLE selects only the top-1 expert for specific tasks.
-In dense activation, MoLE [27] treats each trained LoRA module at each layer as an individual expert, integrating learnable gating functions at each layer to control hierarchical weights.
-In rectified activation, RoDE [10] designs diverse experts and employs a rectified linear unit to optimize the router for sparse expert activation.
-These methods involve complex expert group configurations or routing strategies.
-In contrast, our approach offers a simple yet efficient adapter that effectively addresses data conflicts without requiring a complicated design.
-Figure 3: Illustration of (a) VCE, (b) typical Mixture of Experts strategies for addressing data conflicts, and (c) Dual-LoRA.
-Mixture of Experts strategies include: Sparse activation [4], where only a subset of experts is activated; dense activation [27], where all experts contribute equally; and rectified activation [10], where all experts contribute with selective rectification.
+While most of these methods require full fine-tuning of the LLM, our approach, in contrast, introduces a lightweight Visual Cue enhancement module for efficient fine-tuning.
+Figure 3: Overview of our two key components: Visual Cue Enhancement (VCE) module enhances high-level anchor features by aggregating local information from multi-level feature maps.
+The Dual Low-Rank Adaptation (Dual-LoRA) module projects the input feature into two low-rank subspaces: one for stable holistic domain knowledge (skill space) learning and the other for instruction condition (task space) learning.
+Dual-LoRA modules is integrated into the LLM’s query and value projection layers for efficiency.
 3 Method
@@
 For convenience, we omit the indices $i$ in the following.
-With LoRA, the trainable parameters of $F$ are reduced to $\{\theta_{P},A,B\}$.
+Incorporating LoRA, the trainable parameters of $F$ are reduced to $\{\theta_{P},A,B\}$.
 3.2 Framework Overview
-We propose a two-stage approach for efficient Visual Instruction Tuning: (1) Enhanced Vision Projector Pretraining and (2) Dual Low-Rank Adaptation (Dual-LoRA) Visual Instruction Fine-Tuning, as illustrated in Figure 1.
-Specifically, to enable comprehensive vision information processing and effective execution of instruction-following downstream tasks, we propose two main components: Visual Cue Enhancement (VCE) to capture enriched visual cues by incorporating multi-level visual features, and Dual-LoRA to mitigate data conflicts by decoupling visual instruction tuning into skill-task dual low-rank space learning.
-The complete framework pipeline is illustrated in Figure 2, with further details provided in the following sections.
+A visual overview is provided in Fig.3.
+The proposed architecture introduces two key components: 1) a Visual Cue Enhancement (VCE) module, detailed in Sec.3.3, which enhances vision feature projection by aggregating multi-level local visual features, and 2) Dual Low-Rank Adaptation (Dual-LoRA), described in Sec.
+3.4, which employs dual low-rank adaptation pathways to facilitate holistic-to-localized knowledge adaptation.
 3.3 Multi-level Local Visual Cue Enhancement
-Typical vision projectors primarily rely on high-level vision feature maps to align vision and language modalities.
-For example, LLaVA [14] and Qwen-VL [1] use the penultimate layer output of ViT, which captures high-level, language-aligned semantic features but may overlook finer, detail-oriented visual cues.
-To address these limitations, we propose a lightweight Vision Cues Enhancement (VCE) module.
-This module uses the final vision feature map as an anchor feature and applies a single deformable cross-attention mechanism to neighboring patches across other layer feature maps for each patch of the anchor feature.
-An intuitive illustration is shown in Figure 3 (a).
-Denote $F=[f]$ as the multi-level feature maps produced by the vision encoder $V$.
-We refer to the high-level feature map used for projection into vision tokens as the anchor feature $f^{\text{a}}$, while several intermediate layers are designated as reference feature maps $[f^{\text{ref}}]$.
-The vision encoder processes the input image $I$ through multiple blocks to extract these multi-level features $F$, formulated as:
-[…]
-where select-index specifies the index of reference feature maps.
-To enhance the vision cues, we integrate local visual details from $[f^{\text{ref}}]$ into $f^{\text{a}}$.
-We utilize deformable attention [28] to efficiently compute cross-attention across multi-level feature maps, applied between $[f^{\text{ref}}]$ and $f^{\text{a}}$:
-$f^{*}=\text{deform-attn}(f^{\text{a}},[f^{\text{ref}}];\theta_{\text{deform-attn}}),$ (6)
-[…]
-After obtaining the anchor feature $f^{\text{a}}$ and the enhanced visual cue $f^{*}$, we add them up and apply normalization to produce the final enhanced vision feature map $f_{\text{vision}}$.
-This feature map is subsequently projected into vision tokens $t_{\text{vision}}$ through the vision feature projector $P$:
-$f_{\text{vision}}$ $=\text{norm}(f^{\text{a}}+f^{*}),$ (7)
-$t_{\text{vision}}$ $=P(f_{\text{visual}};\theta_{P}).$ (8)
-The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the large language modal model $G$ for further processing.
-The structure of the adaptive $G$ is detailed in the next section.
+Typical MLLMs rely on high-level vision feature maps to align vision and language modalities, e.g.
+LLaVA [20] and Qwen-VL [3] use the penultimate layer vision feature map of ViT [28].
+While this map primarily captures high-level semantic features and may overlook local visual details.
+To address this limitation, we propose a lightweight Visual Cue Enhancement (VCE) module.
+This module enhances the projected feature map, i.e., the high-level feature $*F^{\text{*}}$, by integrating local information from other layers.
+Let $[F_{l}]_{l=1}^{L}$ denote multi-level vision feature maps extracted from selected intermediate layers, where $l\in\{1,\ldots,L\}$ indexes the select layers.
+Suppose that the anchor patch feature to be enhanced is $F^{*}(p_{q})$, where $p_{q}\in\mathbb{R}^{2}$ is its 2D position.
+For each layer $l$, we employ deformable attention [34, 38] to extract local feature $F_{l}^{\prime}(p_{q})$:
+$F_{l}^{\prime}(p_{q})$ $=\sum_{m=1}^{M}\sum_{k=1}^{K}A_{lmq}(k)\cdot W_{lm}F_{l}\left(p_{q}+\Delta p_{lmq}(k)\right),$ (5)
+$A_{lmq}$ $=\mathrm{softmax}\left(\phi_{lm}^{a}(F_{l}(p_{q}))\right),$ (6)
+$\Delta p_{lmq}$ $=\phi_{lm}^{o}(F_{l}(p_{q})),$ (7)
+where $m\in\{1,\ldots,M\}$ indexes attention heads, $W_{lm}$ is a learnable projection weight for head $m$ at layer $l$, and $k\in\{1,\ldots,K\}$ indexes reference patches.
+The offset $\Delta p_{lmq}\in\mathbb{R}^{K\times 2}$ represents adaptively learned coordinate offsets around $p_{q}$, indicate the interest ares, while $A_{lmq}\in\mathbb{R}^{K}$ denotes attention weights for input $F_{l}(p_{q})$, where $\phi_{lm}^{o}$ and $\phi_{lm}^{a}$ are linear layers used for learning the offsets $\Delta p_{lmq}$ and attention $A_{lmq}$, respectively.
+A Multi-Level Local Information Aggregation module is then used to consolidate local features across layers:
+$F^{\prime}(p_{q})=W_{o}[\mathrm{concat}(F_{l}^{\prime}(p_{q}))_{l=1}^{L}],$ (8)
+where $W_{o}$ projects the concatenated features into a unified space.
+Finally, the enhanced feature map is obtained through residual fusion and sent to the vision projector to generate enhanced vision tokens:
+$t_{\mathrm{vision}}=P\left(\mathrm{Norm}\left(F^{*}+\gamma F^{\prime}\right);\theta_{P}\right),$ (9)
+where $\gamma$ is a scaling parameter, and $\mathrm{Norm}(\cdot)$ denotes a LayerNorm [2] layer applied to normalize the residual.
 3.4 Dual Low Rank Adaptation
-LoRA assumes that in large pre-trained models, the model parameters have a high degree of redundancy, with a low-rank structure that can be effectively represented through low-rank matrix approximation.
-However, as downstream tasks become more complex or diverse, data conflicts can arise due to the limited representational capacity of the low-rank structure [9, 4].
-Some methods integrate Mixture of Experts into LoRA to better adapt to diverse downstream tasks.
-They propose expert modules with different granularities [10] and use various routing strategies [4, 27, 18, 10].
-In contrast, we propose Dual Low-Rank Adaptation (Dual-LoRA), a unified approach that decouples adaptation into a task-skill dual low-rank structure, removing the need for complex expert splitting or intricate routing strategies.
-Figure 3 (b) and (c) illustrate the differences between our method and other strategies for mitigating data conflicts.
+In LoRA-MoE, the token embedding is passed through a router, typically a linear layer, followed by an activation strategy that determines which LoRA experts to activate.
+The activation strategies typically include sparse activation [5], dense activation [33], and rectified activation [13], each of which has distinct effects across various tasks.
+However, efficiently balancing the activation strategy, the granularity of LoRA experts, and task complexity can be challenging.
+Moreover, the data passing through the router and multiple LoRA experts significantly increases inference time.
+For instance, a 4-expert LoRA-MoE uses 1.59 times the inference time compared to the non-MoE method (details can be found in Fig.
+8).
+Hence, we aim to implement a unified approach that achieves the same effect as the LoRA-MoE method but more efficiently.
+We begin by analyzing the expressiveness of both LoRA-MoE and unified LoRA.
 Proposition 1: For a set of LoRAs $\{B_{k}A_{k}\}_{k=1}^{K}$, each with rank 1, where $K$ is the total number of LoRAs, the representation space of a single LoRA with rank $K$ contains, and is at least as expressive as, the union of the representation spaces of the individual rank-1 LoRAs.
 Corollary 1: For a set of LoRAs $\{B_{k}A_{k}\}_{k=1}^{K}$, the combined representation capability does not exceed that of a single LoRA with rank $\sum\text{rank}(\text{LoRA}_{k})$.
+The proof of Proposition 1 and its Corollary 1 is provided in Appendix A.
+3.4.1 Learning Holistic Knowledge in One Skill Space
+According to Corollary 1, a single LoRA with equivalent parameters has the same expressiveness as LoRA experts across various ranks.
+Let $\mathcal{S}$ denote the adaptation space, which we refer to as the skill space, modeled by the parameter matrix $S$.
+In practice, however, LoRA-MoE often outperforms a single, larger LoRA [13].
+What leads to the performance gap, despite they have the same expressiveness?
+This can be attributed to the differentiated local response capability of LoRA-MoE, i.e., expert activation, which provides redundant space to address data conflicts.
+Two methods ensure the local differences among experts are proposed: 1) variations in experts’ weight distributions at a given moment, e.g., RoDE [13] employs heterogeneous LoRA experts initialized with varying distributions, and 2) differences in routing scores among experts, which can lead to uneven expert activation, e.g., in extreme cases, only a single expert may be utilized [5].
+The observation above suggests that $\mathcal{S}$ should respond locally (specifically) to the varying tasks. We now consider how to elegantly rectify $\mathcal{S}$ to achieve a “local response”.
 Corollary 2: For a LoRA $BA$ with rank $r$, the LoRA $B(A\odot\sigma(T))$ can be decomposed into any set of LoRA groups $\{B_{k}A_{k}\}_{k=1}^{K}$, provided that the constraint $r\geq\sum_{k=1}^{K}\text{rank}(B_{k}A_{k})$ holds.
 Here, $T$ is a matrix with the same shape as $A$, and $\sigma$ is a non-linear activation function.
-The proofs of Proposition 1 and its corollary are provided in Supplementary Materials A.
-Inspired by these findings, we aim to explore the feasibility of using a single larger-rank LoRA to achieve the same effect mixture of LoRA experts.
-3.4.1 Skill Low-Rank Space
-We propose the construction of a skill low-rank space to encapsulate specific knowledge for downstream tasks while ensuring domain knowledge consistency.
-According to Corollary 1, a single LoRA with equivalent parameters can theoretically represent any configuration of LoRA experts across various ranks without loss of expressive power.
-Let $\mathcal{S}$ represent the skill space of the unified LoRA, modeled by a parameter metric $S$.
-In practice, however, a split LoRA expert group often outperforms a single, larger LoRA [10].
-This can be attributed to variations in routing strategies and initialization across individual LoRAs.
-For example, LLaVA-MoLE [4] activates the top-k experts for targeted learning, while RoDE [10] employs heterogeneous LoRA experts, each initialized with distinct distributions.
-The complex training process, particularly the routing score, leads to divergence in the representational spaces learned by different LoRAs over time.
-Intuitively, even for the same input, each expert may experience unique gradients due to its intrinsic numerical differences and routing scores.
-This observation suggests the need for an additional space to mediate between the instruction task and the skill space $\mathcal{S}$, allowing $S$ to be allocated with numerical differences according to instruction variance.
-3.4.2 Rectified Skill Low-Rank Space
-Based on the skill low-rank space, we aim to activate prior knowledge in the task space guided by specific instructions.
-According to Corollary 2, additional space can be used to map the skill space $\mathcal{S}$, enabling the simulation of sparsely activated experts that respond to varying instructional tasks.
-We use the parameter matrix $T$ to modulate $S$, which is mapped through a non-linear activation $\sigma$ to achieve the effect of rank rectification.
-In this paper, we use the ReLU [11] activation function due to its sparsity-inducing properties and ease of optimization.
-The rectified, decomposed feature space, mapped from the skill and task spaces, can then be expressed as:
+The proof of Corollary 2 is provided in Appendix A.
+3.4.2 Localized Adaptation by Skill Space Rectification
+According to Corollary 2, additional space can be utilized to map $\mathcal{S}$ to simulate the “local response”, which we refer to as the task space $\mathcal{T}$.
+We use a parameter matrix $T$ to model $\mathcal{T}$, $T$ is mapped through a non-linear activation $\sigma$ to achieve the effect of rank rectification.
+In this paper, we use the ReLU [17] activation function due to its sparsity-inducing properties and ease of optimization [10, 13].
+The rectified space, mapped from the skill and task spaces, can then be expressed as:
 $D(x)$ $=\frac{r}{\alpha}B(Sx\odot\sigma(Tx)),$ (10)
 where $D$ denotes the Dual-LoRA module.
-To smooth the distribution of the skill space and improve stability in subsequent computations, we normalize the skill space using layer normalization, which can be easily implemented thanks to the unified single LoRA design.
+To smooth the distribution of the skill space, and improve stability in subsequent computations, we normalize the skill space using layer normalization [2], which can be easily implemented thanks to the holistic skill space design.
 With normalization applied to the skill space, Equation 10 can be revised as:
@@
 $z$ $=Wx+\frac{r}{\alpha}B(\text{Norm}(Sx)\odot\sigma(Tx)).$ (12)
+Remark: Dual-LoRA has a much simpler structure than LoRA-MoE and offers significantly higher time efficiency, an inference time effective comparison is shown in Fig. 8.
 3.5 Training
-Our method consists of two training stages, as shown in Figure 1.
-In the vision projector pretraining stage, we train the parameters of the vision projection modules, i.e., $\{\theta_{P},\theta_{\text{deform-attn}}\}$, using descriptive image-caption pairs.
-In the subsequent visual instruction fine-tuning stage, we enable additional training for the Dual-LoRA parameters, where $\{\theta_{P},\theta_{\text{deform-attn}},S,T,B\}$ are set to be trainable, leveraging the full range of downstream tasks.
+We employ a two-stage fine-tuning framework, consisting of: Stage 1: Vision Projector Pre-training, where only the VCE module and vision feature projector are trainable; and Stage 2: Visual Instruction Fine-tuning, where the VCE module, the vision projector, and the injected Dual-LoRA module are all set to trainable.
 Method Ingredient Recognition Recipe Generation Nutrition Estimation (pMAE $\downarrow$ )
@@
 vanilla LoRA [11] 23.2 34.1 12.4 40.1 46.2 45.5 57.1 53.4 48.7 50.2
-MoE-LoRA (top-2) [27] 22.9 33.8 12.7 40.2 45.56 45.8 56.9 54.4 48.0 50.1
-MoE-LoRA (softmax) [27] 22.7 33.5 12.5 40.0 45.3 45.5 58.1 53.7 47.5 50.0
+LoRA-MoE (top-2) [33] 22.9 33.8 12.7 40.2 45.56 45.8 56.9 54.4 48.0 50.1
+LoRA-MoE (softmax) [33] 22.7 33.5 12.5 40.0 45.3 45.5 58.1 53.7 47.5 50.0
 RoDE [13] 23.6 34.6 13.8 41.4 45.8 47.6 58.5 54.4 50.4 51.3
@@
 Table 1: Performance comparison with state-of-the-art Mixture of Experts in LoRA approaches on UniFood dataset.
-For fair comparison, all methods use the same base pretrained MLLM, LLaVA-1.5-7B.
+All methods use the same base pretrained MLLM, LLaVA-1.5-7B.
 The best performance is highlighted in bold, and the second-best is underlined.
-[…]
-InstructBLIP [6] Vicuna-7B EVA-CLIP ViT-G 224 36.0 60.9 58.8 - 26.2
-IDEFICS [3] LLaMA-7B OpenCLIP ViT-H/14 224 48.2 - 44.5 - -
-Shikra [3] Vicuna-7B CLIP ViT-L/14 224 58.8 - - - -
-Qwen-VL [1] Qwen-7B OpenCLIP ViT-bigG 448 38.2 - 62.3 - -
-Qwen-VL-Chat [1] Qwen-7B OpenCLIP ViT-bigG 448 60.6 65.4 - - -
-LLaVA-1.5 [15] Vicuna-7B CLIP ViT-L/14 336 64.3 - 65.4 85.9 31.1
-LLaVA-1.5-LoRA [15] Vicuna-7B CLIP ViT-L/14 336 66.1 - 67.9 86.4 30.2
-Dual-LoRA Vicuna-7B CLIP ViT-L/14 336 65.1 64.5 67.4 86.8 31.5
-Dual-LoRA+VCE Vicuna-7B CLIP ViT-L/14 336 65.6 64.7 68.1 87.2 32.1
-Table 2: Performance comparison with state-of-the-art MLLMs on general benchmarks. The best performance is highlighted in bold, and the second-best is underlined.
+Method ScienceQA Flickr30k
+Accuracy $\uparrow$ ScareBLEU $\uparrow$ Rouge-L $\uparrow$
+Vanilla LoRA 70.01 27.89 66.76
+LoRA-MoE (top-2) 76.3 28.15 66.58
+LoRA-MoE (softmax) 77.73 28.06 67.18
+RoDE 78.39 28.17 66.89
+Dual-LoRA 79.17 28.25 67.35
+Dual-LoRA+VCE 80.01 28.71 67.86
+Table 2: Accuracy of single-choice questions on ScienceQA and Image captioning performance on Flickr30k. The best performance is highlighted in bold, and the second-best is underlined.
 4 Experiments
-4.1 Model and Optimizer Configuration
-For the base pre-trained multimodal large language model, we use the widely adopted open-source model LLaVA-1.5-7B [14].
-LLaVA-1.5-7B integrates a CLIP ViT-L [22] with an image resolution of 336px and a patch size of 14, a two-layer MLP projector to map visual features into tokens, and Vicuna v1.5 [5] as the language model.
-In our setup, we omit any special tokens, such as image indicator tokens, to enclose visual tokens.
-For the VCE module configuration, we use the outputs from the 2nd-to-last, 8th-to-last, 14th-to-last, and 20th-to-last layers of CLIP ViT-L as reference feature maps (see Section 3.3 for details).
-In the Dual-LoRA setting, we set the LoRA dropout rate to 0.05.
-Our VCE module requires only 5.52 MB of memory, compared to the 40.02 MB needed for the vision projector in LLaVA-1.5-7B.
-Furthermore, our Dual-LoRA task consumes approximately 1.5 MB per LoRA rank when injected solely into the query projection and value projection layers.
-For optimization, we use the Adam optimizer across all trainable parameters.
-During the pretraining stage, we set a global batch size of 256 and a learning rate of 0.001 with no weight decay.
-In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LLM adapter and 0.00002 for the vision feature projector, respectively.
-4.2 Evaluation on Downstream Tasks
-4.2.1 Experiment Setup
-Dataset and Downstream Tasks.
-We begin by evaluating our proposed method on downstream fine-tuning tasks using the UniFood dataset [10], a multi-task dataset with 100,000 samples in the food domain.
-The tasks in UniFood include ingredient recognition, recipe generation, and nutrition estimation.
+4.1 Experimental setup
+Datasets and evaluation metrics.
+We begin by evaluating our proposed method on downstream tasks using the following datasets and metrics: 1) UniFood [13] is a food-related dataset that includes: ingredient recognition, recipe generation, and nutrition estimation tasks.
 For ingredient recognition, we use Intersection over Union (IoU) and F1-score as evaluation metrics.
-For recipe generation, we assess performance with SacreBLEU [21] and Rouge-L [13].
-For nutrition estimation, we use the percentage of mean average error (pMAE) [32] as the metric.
-Baselines.
-We select several Mixture of Experts (MoE) methods within the LoRA framework: a sparse activation method, MoE-LoRA (top-2), which activates the top-2 experts; a dense activation method, MoE-LoRA (softmax), which uses a softmax router to engage all experts; and a rectified activation method, RoDE [10], with rectified diverse expert activation.
-Both MoE-LoRA (top-2) and MoE-LoRA (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$.
-Additionally, we include a vanilla LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
+For recipe generation, we assess performance using SacreBLEU [27] and Rouge-L [19].
+For nutrition estimation, we use the percentage of mean average error (pMAE) [36] as the evaluation metric.
+2) Flickr30k [26] is a large image description dataset, with each sample having 5 descriptions.
+All 5 descriptions are used for training to simulate potential data conflicts, SacreBLEU and Rouge-L metrics are used for evaluation.
+3) ScienceQA [29] is a large-scale multimodal science question-answering benchmark covering diverse subjects.
+We evaluate it using single- choice questions, with average accuracy as the metric.
+All datasets are trained for one epoch on the training set and evaluated on the test set.
+Baseline methods.
+We select several Mixtures of LoRA Experts (LoRA-MoE) methods within the LoRA framework: a sparse activation method, LoRA-MoE (top-2) [33], which activates the top-2 experts; a dense activation method, LoRA-MoE (softmax) [33], which uses a softmax router to engage all experts; and a rectified activation method, RoDE [13], with rectified diverse expert activation.
+Both LoRA-MoE (top-2) and LoRA-MoE (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$.
+Additionally, we include a standard LoRA baseline with a rank of 64, and our Dual-LoRA is also set to rank 64.
 Implementation Details.
+We use the widely adopted open-source model LLaVA-1.5-7B [20] for the base MLLM.
 We conduct our experiments on a setup with 8 RTX 4090 GPUs.
-The LLM adapters are integrated only within the linear layers of the query and value projectors in the LLM transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.
-4.2.2 Performance Comparison
-Table 1 presents the results of our method compared to baseline methods on the UniFood dataset.
-As shown, our Dual-LoRA method consistently outperforms the sparse expert activation method (Top-2), the dense expert activation method (Softmax), and the rectified expert activation method (RoDE) across all tasks.
-Additionally, incorporating the proposed VCE module with Dual-LoRA (i.e., Dual-LoRA+VCE) manages to further boost performance for both ingredient recognition and nutrition estimation tasks.
-These results demonstrate the effectiveness of our proposed Dual-LoRA and VCE components for instruction fine-tuning in downstream tasks.
-Figure 4: Qualitative results. Dual-LoRA + VCE produces more accurate and knowledge-consistent answers, whereas the LLaVA-1.5-7b-LoRA generates knowledge-inconsistent responses.
-Figure 5: Performance vs.
-trainable LLM adapter parameters for vanilla LoRA and our proposed Dual-LoRA on the UniFood dataset.
-The adapters are integrated into the query and value projectors of the transformer blocks.
-To eliminate the influence of trainable parameters beyond the LLM adapter, only the LLM adapter is set as trainable.
-4.3 Evaluation on General MLLM Benchmarks
-4.3.1 Experiment Setup
-MLLM Benchmarks.
-For the general benchmark evaluation, following LLaVA [14], we use a 558K subset of the LAION-CC-SBU dataset [23] and the llava-v1.5-mix665k dataset [14] for instruction fine-tuning.
-Four widely used MLLM benchmarks are used for evaluation, including MMBench [17], SEED-Bench [12], and LLaVA-Bench In-the-${}^{\text{W}}$ ) [14].
-The first three benchmarks assess various MLLM capabilities, such as perceptual understanding and visual reasoning, through binary yes/no questions (MME) or multiple-choice questions (MMBench and SEED-Bench).
-We use the image-only subset of SEED-Bench (SEEDI), a popular choice for many image-based MLLMs [2, 30].
-Implementation Details.
-Our experiments are conducted on a setup with 4 $\times$ A100 GPUs (80GB).
-The selected layers for LLM adapter injection match those used in LLaVA-1.5-7B-LoRA [14], with the adapter rank $r$ set to 128 and scaling factor $\alpha$ set to 256 in Equation 11.
-4.3.2 Performance Comparison
-The results of our proposed method, alongside state-of-the-art MLLMs, are presented in Table 2.
-The Dual-LoRA + VCE model consistently ranks among the top two across all benchmark tests.
-Compared to other methods, this model demonstrates robust performance across multiple benchmarks, indicating excellent generalization and strong adaptability to a variety of tasks.
-These findings validate the effectiveness of the VCE module and Dual-LoRA, enabling the model to achieve superior results on general visual understanding tasks.
+The LLM adapters are integrated only within the linear layers of the query and value projectors in the transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.
+The detailed base model and optimization configuration can be found in Appendix B.
+Hyperparameters setting.
+For the VCE module configuration, we select the 2nd, 8th, 14th, and 20th to the last layers of CLIP ViT-L as reference feature maps, with the 2nd layer serving as the anchor feature (see Section 3.3 for details).
+We set the scaling parameter $\gamma$ to 1.0 for all tasks.
+The VCE module requires only 5.52 MB of memory.
+In the Dual-LoRA setting, we set $\alpha$ to 2× the rank and the LoRA dropout rate to 0.05.
 VCE Dual-LoRA Ingredient Recognition Recipe Generation
@@
 The color $\blacksquare$ represents a frozen vision projector, whereas $\blacksquare$ indicates it is trainable.
-Rank LMM Adapter LMM Adapter Memory Size Ingredient Recognition Recipe Generation
-IoU F1 SacreBLEU Rouge-L
-var1 32 LoRA 32M 19.2 29.7 10.4 41.1
-var2 Dual-LoRA 48M 22.7 +18.1% 33.6 +13.3% 12.9 +24.1% 41.3 +0.3%
-var3 64 LoRA 64M 20.6 31.3 12.7 41.3
-var4 Dual-LoRA 96M 23.0 +11.7% 33.7 +7.7% 14.4 +13.1% 41.7 +0.9%
-var5 128 LoRA 128M 21.6 32.5 13.0 41.3
-var6 Dual-LoRA 192M 23.6 +8.9% 34.4 +5.9% 14.8 +14.3% 42.9 +3.6%
-Table 5: Performance and memory usage of the LoRA adapter under varying rank settings on the UniFood dataset. To prevent the effect of vision projector, only the LoRA adapter is set to be trainable.
-4.4 Qualitative Results
-Figure 4 presents a comparison between our proposed method and the vanilla LoRA fine-tuned version of LLaVA-1.5-7B on food-related downstream tasks.
-The vanilla LoRA demonstrates knowledge inconsistency, with the ingredient list failing to align with the generated cooking instructions.
-In contrast, our method produces more accurate and knowledge-consistent answers, ensuring the ingredient list aligns with the cooking instructions.
-These results visually underscore the effectiveness of our approach.
+4.2 Experiment Results
+Performance on Multi-task Tuning.
+Considering the data conflicts induced by inconsistencies in optimization goals during multi-task fine-tuning, we first evaluate our approach on the multi-task datasets UniFood and ScienceQA.
+As shown in Table 1 for the UniFood dataset, our Dual-LoRA method consistently outperforms both vanilla LoRA and LoRA-MoE across all tasks.
+Similarly, for ScienceQA, as illustrated in Fig 2, our method surpasses all baseline approaches.
+The qualitative results in Fig.
+4 further highlight the effectiveness of our approach, as it consistently generates ingredient lists across different tasks, demonstrating strong alignment with holistic domain knowledge.
+Performance on Diverse Annotation Tuning.
+Data conflicts can also arise from diverse annotations within a single task for the same image (e.g.
+varying perspective descriptions of the same image), potentially distorting semantic consistency.
+We use Flickr30k to test our method and baseline approaches in addressing annotation diversity-induced data conflicts.
+As shown in Table 2, our method consistently outperforms the baseline approaches on both SacreBLEU and ROUGE-L scores, further demonstrating the effectiveness of our proposed approach.
+We conduct a more in-depth analysis in the following sections.
+Figure 4: Conversations involving two distinct tasks can lead to knowledge-inconsistent responses with LLaVA-1.5-7b-LoRA, due to conflicts arising from multiple complex tasks.
+In contrast, our proposed method generates more accurate and knowledge-consistent answers.
 4.3 Ablation Study
-In this section, we conduct ablation studies to verify different components of our method on UniFood dataset. The setup is the same as Section 4.2.1.
-VCE and Dual-LoRA.
-We conducted an ablation study to analyze the impact of different projector and adapter configurations on the performance of our model in ingredient recognition and recipe generation.
-The results are shown in Table 3.
-This ablation study shows that both multi-level local visual feature enhancement (VCE) and Dual-LoRA positively impact performance.
-When both are enabled, the best results are achieved for ingredient recognition and recipe generation tasks, indicating that the combination of these methods is more effective than using either one alone.
+Ablation on VCE and Dual-LoRA.
+We conducted an ablation study to evaluate the impact of two key components: the VCE module and the Dual-LoRA module on the UniFood dataset, using the ingredient recognition and recipe generation tasks.
+The results are presented in Table 3.
+This study demonstrates that both the VCE module and the Dual-LoRA module positively influence performance.
+When both modules are enabled, the best results are achieved for both ingredient recognition and recipe generation tasks.
 Dual-LoRA design.
@@
 When both are removed, the performance of Dual-LoRA reaches its lowest level.
+Figure 5: Feature Map Visualization of Enhanced Visual Cue with VCE: (a) Enhanced Visual Cue emphasize key areas in food imagery, such as textures, garnishes, and toppings.
+(b) The cues highlight important details: text for readability, the cheetah-prey interaction, and the woman’s face and Oscar trophy on stage.
+Figure 6: Panels (a), (b), and (c) represent feature visualizations for the recipe generation task where: (a) the distributions of feature outputs for the holistic skill space, (b) the distributions for the rectified skill space, i.e., $\mathcal{S}\odot\mathcal{T}$, and (c) the entropy of the holistic skill space (blue line) and the rectified skill space (orange line).
+Panels (d), (e), and (f) show the corresponding results for the nutrition estimation task.
+Figure 7: Performance vs.
+Trainable LLM Adapter Parameters for Vanilla LoRA and Our Proposed Dual-LoRA on the UniFood Dataset.
+The adapters are integrated into the query and value projectors of the transformer blocks.
+To isolate the effect of trainable parameters beyond the LLM adapters, only the adapters are set as trainable.
+Figure 8: Relative Inference Time Efficiency of Our Dual-LoRA Method and Several LoRA-MoE Methods.
+The inference time of vanilla LoRA is used as the baseline (unit 1).
+The adaptor is injected into the query and value projection layers of the LLM to improve efficiency.
+Test on the Flickr30k test set.
+[…]
+InstructBLIP [7] Vicuna-7B EVA-CLIP ViT-G 224 36.0 60.9 58.8 - 26.2
+Shikra [4] Vicuna-7B CLIP ViT-L/14 224 58.8 - - - -
+LLaVA-1.5 [21] Vicuna-7B CLIP ViT-L/14 336 64.3 - 65.4 85.9 31.1
+LLaVA-1.5-LoRA [21] Vicuna-7B CLIP ViT-L/14 336 66.1 - 67.9 86.4 30.2
+Dual-LoRA Vicuna-7B CLIP ViT-L/14 336 65.1 64.5 67.4 86.8 31.5
+Dual-LoRA+VCE Vicuna-7B CLIP ViT-L/14 336 65.6 64.7 68.1 87.2 32.1
+Table 5: Performance comparison between baseline methods and our proposed method on general benchmarks. The best performance is highlighted in bold, and the second-best in underlined.
+4.4 Qualitative Analysis
+Qualitative Results.
+Fig.
+4 presents a comparison between the vanilla LoRA method and our proposed approach.
+The vanilla LoRA method exhibits knowledge inconsistency, with the ingredient list failing to align with the generated cooking instructions.
+In contrast, our method produces more accurate and knowledge-consistent results, where the ingredient list aligns with the cooking instructions.
+More qualitative results can be found in Appendix D.
+Qualitative Analysis on VCE.
+Fig.
+5 shows the heatmap of the enhanced Visual Cue, i.e., $F^{\prime}$ (see Sec.
+3.3 for details).
+The highlighted areas closely match the regions of interest in the image, demonstrating the effectiveness of our VCE module in enhancing local visual cues.
+Qualitative Analysis of Dual-LoRA.
+Figure 6 offers a comprehensive visualization of both the skill space and it’s rectified space, highlighting several important characteristics: 1) The information entropy of the rectified skill space is lower than that of the original skill space, as shown in panels (c) and (f) of Fig.6, suggesting that the rectified skill space is more information-determined to the current task.
+2) The high-energy frequency bands in the rectified skill space exhibit regional patterns, as evidenced by the varying high values in Fig.6 (b) and (c).
+These observations demonstrate that the rectified skill space presents a more localized activation of the knowledge, emphasizing task-specific features.
 4.5 Efficiency Analysis
-As listed in Table 5, we present the results of the vanilla LoRA and Dual-LoRA across three configurations of total rank values (32, 64, and 128).
-The results indicate that Dual-LoRA consistently outperforms the vanilla LoRA counterparts across all rank settings (32, 64, and 128) with only a slightly higher number of parameters.
-Notably, Dual-LoRA achieves superior performance even with fewer trainable parameters than LoRA with higher ranks, such as Var2-Dual-LoRA vs.
-Var3-LoRA and Var5-LoRA, as well as Var4-Dual-LoRA vs.
-Var5-LoRA, which is also visualized in Figure 5.
-We also observe that Dual-LoRA shows relatively larger improvements over vanilla LoRA when the rank size is small.
-We attribute this to data conflicts being more obvious with fewer trainable parameters, and the incorporation of task space in Dual-LoRA helps mitigate this issue.
-These findings highlight the effectiveness of Dual-LoRA in addressing data conflicts when training multiple tasks in downstream applications.
+Memory Usage Efficiency.
+Fig.
+7 presents the results of vanilla LoRA and Dual-LoRA across three rank configurations (32, 64, and 128).
+As shown, Dual-LoRA consistently outperforms the vanilla LoRA method across all rank settings and tasks.
+Notably, the performance gap is more pronounced with fewer parameters.
+This can be attributed to the increased visibility of data conflicts when fewer parameters are trainable.
+Detailed data on memory usage efficiency comparisons can be found in Appendix E.
+Inference Time Efficiency.
+Fig.
+8 shows the inference time efficiency of our proposed method compared to LoRA-MoE (top-2) [33], LoRA-MoE (softmax) [33], and RoDE [13], with vanilla LoRA as the baseline (unit 1).
+The experimental results reveal that LoRA-MoE implementations exhibit significantly prolonged inference durations, with 2-expert and 4-expert configurations requiring 1.38× and 1.59× baseline latency respectively.
+In contrast, our approach demonstrates exceptional efficiency improvements: the Dual-LoRA configuration achieves only a 1.12$\times$ baseline computational overhead, and when combined with VCE, the complete system maintains high efficiency with merely a 1.16$\times$ increase in baseline latency.
+4.6 Disscusion on General Benchmarks
+We further conduct experiments on general benchmarks for additional analysis.
+The results of our proposed method, alongside state-of-the-art MLLMs, are presented in Table 5, with the experimental setup for these benchmarks detailed in Appendix B.
+The Dual-LoRA + VCE model consistently ranks among the top two across all benchmark tests and achieves the best performance on LLAVAW, POPE and MMVet, demonstrating the strong competitiveness of our method on general tasks.
 5 Conclusion
-In this paper, we explore efficient visual instruction fine-tuning for multi-modal large language models (MLLMs) by leveraging pre-trained models and utilizing only a few adapter parameters.
-Visual instruction fine-tuning of MLLMs faces two key challenges: insufficient capture of visual cues and conflicts among various instruction tasks due to the limited adapter parameters.
-To address these challenges, we propose a framework comprising Vision Cues Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA).
-For Vision cue enhancement, we use multi-level feature maps to enrich the extracted vision cues.
-For Dual Low-Rank Adaptation, we introduce a rectified skill-space mapping technique, enabling the model to learn complex knowledge and respond specifically to instructions.
-Dual Low-Rank Adaptation possesses the expressive power of any ensemble of experts with a combined rank lower than its own.
-Evaluations on both general visual instruction benchmarks and domain-specific benchmarks demonstrate the effectiveness of our proposed method.
+In this paper, we explore the transition from holistic to localized adaptation for efficient visual instruction fine-tuning.
+We introduce Dual Low-Rank Adaptation (Dual-LoRA), which enables the model to capture holistic knowledge while being locally activated for specific tasks, thereby mitigating data conflicts.
+Additionally, we propose Visual Cue Enhancement (VCE), which improves the vision feature projection by incorporating local information from multi-level vision feature maps.
+Our methods are both time- and memory-efficient, with only a 1.16$\times$ increase in time consumption compared to standard LoRA, yet significantly outperform LoRA in terms of performance with the same parameter settings.
+Extensive experiments on downstream tasks validate the effectiveness of our approach.
+Acknowledgements This work is supported by the Science and Technology Commission of Shanghai Municipality (Grant No.
+24511103100).
+Additionally, it is partially supported by the Ministry of Education, Singapore, under the Academic Research Fund (AcRF) Tier 2 (Proposal ID: T2EP20222-0046).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.